Random samples of size 2 are taken from a finite population which consist of measurements 6,7,8,9,10 and 11. Show that the mean of the population is 2.5 and its standard deviation is 0.493
1
Expert's answer
2021-07-13T12:52:53-0400

"N=6 \\

n=2 \\

Number \;of \; samples = \frac{6!}{2!(6-2)!}= 15"



The mean of the population

"\mu = \frac{6+7+8+9+10+ 11}{6}=8.5"

The standard deviation

"\sigma = \sqrt{ \frac{\sum (x- \mu)^2}{n} } \\

\sigma = \sqrt{ \frac{1}{6} ( (6-8.5)^2+(7-8.5)^2+(8-8.5)^2+(9-8.5)^2+(10-8.5)^2+(11-8.5)^2 ) } = 1.70"